# AVSegFormer: Audio-Visual Segmentation with Transformer

### v1

@@ 3.6. Loss Function
| LVS [7] | ResNet-50 [20] | 51.0 | 37.94 | 33.0 | 29.45 | CVPR’2021 | ||
| MSSL [39] | ResNet-18 [20] | 66.3 | 44.89 | 36.3 | 26.13 | ECCV’2020 | ||
| 3DC [36] | ResNet-34 [20] | 75.9 | 57.10 | 50.3 | 36.92 | 21.6 | 17.27 | BMVC’2020 |
| SST [17] | ResNet-101 [20] | 80.1 | 66.29 | 57.2 | 42.57 | CVPR’2021 | ||
| AOT [53] | Swin-B [33] | 31.0 | 25.40 | NeurIPS’2021 | ||||
| iGAN [37] | Swin-T [33] | 77.8 | 61.59 | 54.4 | 42.89 | ArXiv’2022 | ||
| LGVT [55] | Swin-T [33] | 87.3 | 74.94 | 59.3 | 40.71 | NeurIPS’2021 | ||
| AVSBench-R50 [56] | ResNet-50 [20] | 84.8 | 72.79 | 57.8 | 47.88 | 25.2 | 20.18 | ECCV’2022 |
| AVSegFormer-R50 (ours) | ResNet-50 [20] | 85.9 | 76.45 | 62.8 | 49.53 | 29.3 | 24.93 | |
| AVSBench-PVTv2 [56] | PVTv2 [49] | 87.9 | 78.74 | 64.5 | 54.00 | 35.2 | 29.77 | ECCV’2022 |
| AVSegFormer-PVTv2 (ours) | PVTv2 [49] | 89.9 | 82.06 | 69.3 | 58.36 | 42.0 | 36.66 | |
@@
| 81.4 | 89.0 | 57.4 | 67.9 | |
| CRA | 81.7 | 89.7 | 57.8 | 68.2 |
| CHA | 82.0 | 89.9 | 58.3 | 69.3 |
| mask | S4 | MS3 | ||
| reformer | mIoU | F | mIoU | F |
| 81.1 | 88.4 | 57.3 | 68.3 | |
| AR | 81.7 | 89.6 | 57.9 | 68.8 |
| QR | 82.0 | 89.9 | 58.3 | 69.3 |
4. Experiments
4.1. Dataset
@@
AVSBench-Object [57] is an audio-visual dataset specifically designed for the audio-visual segmentation task, containing pixel-level annotations. The videos are downloaded from YouTube and cropped to 5 seconds, with one frame per second extracted for segmentation. The dataset includes two subsets: a semi-supervised single sound source subset for single sound source segmentation (S4), and a fully supervised multi-source subset for multiple sound source segmentation (MS3). S4 subset: The S4 subset contains a total of videos, with videos for training, for validation, and for testing. The target objects cover 23 different categories, including humans, animals, vehicles, and musical instruments. Besides, this subset is trained in a semi-supervised manner, where each video contains five frames, but only the first frame is annotated. MS3 subset: The MS3 subset includes 424 videos, with 286 training, 64 validation, and 64 testing videos, covering the same categories as the S4 subset. Additionally, unlike S4, the MS3 subset is full-supervised with all five frames annotated in training.
@@
AVSBench-Semantic [56] is an extension of the AVSBench-Object dataset, which offers additional semantic labels that are not available in the original AVSBench-Object dataset. It is designed for audio-visual semantic segmentation (AVSS). In addition, the videos in AVSBench-Semantic are longer, with a duration of 10 seconds, and 10 frames are extracted from each video for prediction. It combines semi-supervised and fully-supervised manners, with labels provided for the first and first five frames of videos inherited from the S4 and MS3 subsets, respectively. For the newly collected 7,000 videos, a complete 10-frame label is provided. Overall, the AVSBench-Semantic dataset has increased in size by approximately three times compared to the original AVSBench-Object dataset, with 8,498 training, 1,304 validation, and 1,554 test videos.
Metric. These benchmark datasets employ mean intersection over union (mIoU) and F-score as the evaluation metrics.
4.2. Implementation Details
Our method is evaluated on three AVS sub-tasks, including single sound source segmentation (S4), multiple sound source segmentation (MS3), and audio-visual semantic segmentation (AVSS).
@@ 4.2. Implementation Details
Details. We train our AVSegFormer models for the three sub-tasks using an NVIDIA V100 GPU. Consistent with previous works [57, 56], we employ AdamW [35] as the optimizer, with a batch size of 4 and an initial learning rate of . All video frames are resized to 224 224 resolution. For the S4 and MS3 subsets, each video contains 5 frames, while each video in AVSS contains 10 frames. Since the MS3 subset is quite small, we train it for 60 epochs, while the S4 and AVSS subsets are trained for 30 epochs. The encoder and decoder in our AVSegFormer both are comprised of 6 layers with an embedding dimension of 256. We set the coefficient of the proposed intermediate mask loss to 0.1 for the best performance. More detailed training settings can be found in Table 3.
4.3. Comparison with Prior Arts
@@
To verify the effectiveness of our method, we conducted a comprehensive comparison between our AVSegFormer and existing methods on the AVS benchmark [57, 56]. For fairness, we employ the ImageNet-1K [14] pre-trained ResNet-50 [20] or PVTv2 [49] as the backbone to extract visual features, and the AudioSet [18] pre-trained VGGish [21] to extract audio features.
@@
Comparison with methods from related tasks. Firstly, we compare our AVSegFormer with state-of-the-art methods from three AVS-related tasks, including sound source localization (LVS [7] and MSSL [39]), video object segmentation (3DC [36], SST [17] and AOT [53]), and salient object detection (iGAN [37] and LGVT [55]). These results are collected from the AVS benchmark [57, 56], which are transferred from the original tasks to the AVS tasks.
@@
As shown in Table 1, our AVSegFormer exceeds these methods by large margins. For instance, on the S4 subset, AVSegFormer-R50 achieves an impressive mIoU of 76.45, which is 1.51 points higher than the best LGVT. Although LGVT has a better Swin-T [33] backbone, our AVSegFormer with ResNet-50 backbone still performs better regarding mIoU. In addition, AVSegFormer-PVTv2 produces an outstanding mIoU of 82.06 and an F-score of 89.9 on this subset, which is 7.12 mIoU and 2.6 F-score higher than LGVT, respectively. On the MS3 subset, AVSegFormer-R50 outperforms the best iGAN with 6.64 mIoU and 8.4 F-score, while AVSegFormer-PVTv2 further raised the bar with an exceptional improvement of 15.47 mIoU and 14.9 F-score. On the AVSS subset, our AVSegFormer-R50 yields 24.93 mIoU and 29.3 F-score, which are slightly lower than the best method AOT. Nevertheless, AVSegFormer-PVTv2 obtains an impressive performance of 36.66 mIoU and 42.0 F-score, surpassing AOT by 11.26 mIoU and 11.0 F-score, respectively.
@@
Comparison with AVSBench baseline [57, 56]. Then, we compare our AVSegFormer with the AVSBench baseline, which is the current state-of-the-art method for audio-visual segmentation. As reported in Table 1, on the S4 subset, AVSegFormer-R50 achieves 3.66 mIoU and 1.1 F-score improvements over AVSBench-R50, while AVSegFormer-PVTv2 surpasses AVSBench-PVTv2 by 3.32 mIoU and 2.0 F-score. On the MS3 subset, AVSegFormer-PVTv2 surpasses AVSBench-PVTv2 with a margin of 4.36 mIoU and 4.8 F-score. On the AVSS subset, AVSegFormer-R50 and AVSegFormer-PVTv2 achieve significant results with an mIoU improvement of 4.75 and 6.89, and a substantial F-score improvement of 4.1 and 6.8. These results demonstrate that AVSegFormer outperforms the AVSBench baseline on all sub-tasks, becoming a new state-of-the-art method for audio-visual segmentation.
4.4. Ablation Study
@@
In this section, we conduct ablation experiments to verify the effectiveness of each key design in the proposed AVSegFormer. Specifically, we adopt PVTv2 [49] as the backbone and conduct extensive experiments on the S4 and MS3 sub-tasks. Other training settings are the same as in Section 4.2.
@@
Effect of learnable queries. We further investigated the impact of learnable queries in the decoder inputs. As shown in Table 2b, the improvement due to the learnable queries is relatively small in the single sound source task (S4), while it brings significant improvement in the multiple sound source task (MS3). This can be attributed to the complexity of sounding objects. In S4, since there is only one sounding object that remains unchanged, its auditory features are distinct enough. In contrast, in MS3, it becomes relatively difficult for the model to learn target features solely based on audio features, due to the mixing and change of sound sources. The learnable queries partially compensate for this limitation. The results indicate that although the concept of learnable queries is simple, it can produce excellent results in complex scenarios.
@@
Effect of audio-visual mixer. We then studied the impact of the audio-visual mixer in our model. There are two versions designed for this module, as illustrated in Figure 4. The cross-attention mixer (CRA) utilizes visual features as queries and audio features as keys/values for cross-attention, aiming to bring audio information into visual features in the early stages. The channel-attention mixer (CHA) introduced the mechanism of channel attention with audio features as queries and visual features as keys/values. As presented in Table 2c, the design of CHA brought greater performance improvement compared to CRA.
@@
In addition, to validate our suppose that the audio-visual mixer can amplify relevant features while suppressing irrelevant ones, we also visualize features before and after the audio-visual mixer along with their predicted masks. Figure 6 presents the visualization results. We compare the original features with the mixed features generated by the audio-visual mixer and selected 9 channels for rendering. It is evident that for the sounding object (right girl and dog), the mixer effectively enhanced its features. Meanwhile, the non-sounding objects (left girl, table/bed, or background) experienced some degree of suppression. As a result, the predicted mask without mixer may segment out the wrong target (left girl), and the mixer can lead to a more accurate prediction of the correct sounding object. These findings align with our hypothesis and further substantiate the effectiveness of the audio-visual mixer.
Effect of intermediate mask loss. We finally conducted experiments to learn the impact of the intermediate mask loss. Similarly, we designed two versions for the mask reformer module as shown in Figure 5. In the attention-only reformer (AR), we directly feed the attention map of each cross-attention layer into an FPN network to generate intermediate masks. In the query-mixed reformer (QR), we multiply the attention map with the output queries of the corresponding decoder layer, and then feed the generated features into the FPN. The experimental results presented in Table 2d demonstrate that both versions have brought some performance improvement, with the adopted QR outperforming AR.
@@
Qualitative analysis. We also present the visualization results of AVSegFormer compared with those of AVSBench on three audio-visual segmentation tasks in Figure 7. The top line of each task displays the raw images, and the second line displays the ground truth. The last two lines display the predicted masks by AVSBench and AVSegFormer, respectively. The visualization results clearly demonstrate that our method performs better than the previous method. It has a strong ability in target localization and semantic understanding, which can accurately segment and classify the sounding objects in each task. Furthermore, in multiple sound sources scenes, the model can effectively identify the correct sound source and accurately segment the target object. These results highlight the effectiveness and robustness of our method.
5. Conclusion
In this paper, we propose AVSegFormer, a novel framework that leverages the power of transformers to achieve leading performance in audio-visual segmentation tasks. First, our method introduces learnable and audio queries, enabling our network to dynamically attend to relevant visual features and significantly enhance segmentation performance. Second, we design an audio-visual mixer that selectively amplifies or suppresses different visual channels, making the visual features better adapt to diverse audio inputs. Additionally, we propose an intermediate mask loss to improve the effectiveness of the training process. Our experimental results demonstrate the superior performance of AVSegFormer compared to existing state-of-the-art methods, and a series of ablation studies validate the effectiveness of our proposed components.
Appendix A Appendix
We listed the detailed settings of our models in Table 3.
| Settings | S4 | MS3 | AVSS |
| input resolution | |||
| frames | 5 | 5 | 10 |
| embedding dimension | 256 | 256 | 256 |
| transformer layers | 6 | 6 | 6 |
| number of queries | 300 | 300 | 300 |
| mask loss coefficient | 0.1 | 0.1 | 0.1 |
| learnable queries | ✓ | ✓ | ✓ |
| audio-visual mixer | ✓ | ✓ | ✓ |
| batch size | 4 | 4 | 4 |
| optimizer | AdamW | AdamW | AdamW |
| learning rate | |||
| drop path rate | 0.1 | 0.1 | 0.1 |
| epoch | 30 | 60 | 30 |

### v4

@@ Discussion
LVS | ResNet-50 | 51.0 | 37.94 | 33.0 | 29.45 | CVPR’2021 | ||
MSSL | ResNet-18 | 66.3 | 44.89 | 36.3 | 26.13 | ECCV’2020 | ||
3DC | ResNet-34 | 75.9 | 57.10 | 50.3 | 36.92 | 21.6 | 17.27 | BMVC’2020 |
SST | ResNet-101 | 80.1 | 66.29 | 57.2 | 42.57 | CVPR’2021 | ||
AOT | Swin-B | 31.0 | 25.40 | NeurIPS’2021 | ||||
iGAN | Swin-T | 77.8 | 61.59 | 54.4 | 42.89 | ArXiv’2022 | ||
LGVT | Swin-T | 87.3 | 74.94 | 59.3 | 40.71 | NeurIPS’2021 | ||
AVSBench-R50 | ResNet-50 | 84.8 | 72.79 | 57.8 | 47.88 | 25.2 | 20.18 | ECCV’2022 |
DiffusionAVS-R50 | ResNet-50 | 86.9 | 75.80 | 62.1 | 49.77 | ArXiv’2023 | ||
AVSegFormer-R50 (ours) | ResNet-50 | 85.9 | 76.45 | 62.8 | 49.53 | 29.3 | 24.93 | AAAI’2024 |
AVSegFormer-R50* (ours) | ResNet-50 | 86.7 | 76.38 | 65.6 | 53.81 | 31.5 | 26.58 | AAAI’2024 |
AVSBench-PVTv2 | PVTv2 | 87.9 | 78.74 | 64.5 | 54.00 | 35.2 | 29.77 | ECCV’2022 |
DiffusionAVS-PVTv2 | PVTv2 | 90.2 | 81.38 | 70.9 | 58.18 | ArXiv’2023 | ||
AVSegFormer-PVTv2 (ours) | PVTv2 | 89.9 | 82.06 | 69.3 | 58.36 | 42.0 | 36.66 | AAAI’2024 |
AVSegFormer-PVTv2* (ours) | PVTv2 | 90.5 | 83.06 | 73.0 | 61.33 | 42.8 | 37.31 | AAAI’2024 |
@@
81.4 | 88.0 | 59.4 | 70.9 | |
CRA | 82.7 | 89.7 | 59.8 | 72.2 |
CHA | 83.1 | 90.5 | 61.3 | 73.0 |
mix | S4 | MS3 | ||
loss | mIoU | F | mIoU | F |
✓ | 83.1 | 90.5 | 61.3 | 73.0 |
81.3 | 89.1 | 59.6 | 71.3 |
Dataset
@@
AVSBench-Object (Zhou et al. 2022) is an audio-visual dataset specifically designed for the audio-visual segmentation task, containing pixel-level annotations. The videos are downloaded from YouTube and cropped to 5 seconds, with one frame per second extracted for segmentation. The dataset includes two subsets: a single sound source subset for single sound source segmentation (S4), and a multi-source subset for multiple sound source segmentation (MS3). S4 subset: The S4 subset contains videos, with videos for training, for validation, and for testing. The target objects cover 23 categories, including humans, animals, vehicles, and musical instruments. MS3 subset: The MS3 subset includes 424 videos, with 286 training, 64 validation, and 64 testing videos, covering the same categories as the S4 subset.
@@
AVSBench-Semantic (Zhou et al. 2023) is an extension of the AVSBench-Object, which offers additional semantic labels that are not available in the original AVSBench-Object dataset. It is designed for audio-visual semantic segmentation (AVSS). In addition, the videos in AVSBench-Semantic are longer, with a duration of 10 seconds, and 10 frames are extracted from each video for prediction. Overall, the AVSBench-Semantic dataset has increased in size by approximately three times compared to the original AVSBench-Object dataset, with 8,498 training, 1,304 validation, and 1,554 test videos.
Implementation Details
@@
We train our AVSegFormer models for the three AVS sub-tasks using an NVIDIA V100 GPU. Consistent with previous works (Zhou et al. 2022, 2023), we employ AdamW (Loshchilov and Hutter 2017) as the optimizer, with a batch size of 2 and an initial learning rate of . Since the MS3 subset is quite small, we train it for 60 epochs, while the S4 and AVSS subsets are trained for 30 epochs. The encoder and decoder in our AVSegFormer comprise 6 layers with an embedding dimension of 256. We set the coefficient of the proposed mixing loss to 0.1 for the best performance. More detailed training settings can be found in the supplementary materials.
Comparison with Prior Arts
@@
We conducted a comprehensive comparison between our AVSegFormer and existing methods (Zhou et al. 2022, 2023; Mao et al. 2023) on the AVS benchmark. For fairness, we employ the ImageNet-1K (Deng et al. 2009) pre-trained ResNet-50 (He et al. 2016) or PVTv2 (Wang et al. 2022b) as the backbone to extract visual features, and the AudioSet (Gemmeke et al. 2017) pre-trained VGGish (Hershey et al. 2017) to extract audio features.
@@
Comparison with methods from related tasks. Firstly, we compare our AVSegFormer with state-of-the-art methods from three AVS-related tasks, including sound source localization (LVS (Chen et al. 2021) and MSSL (Qian et al. 2020)), video object segmentation (3DC (Mahadevan et al. 2020), SST (Duke et al. 2021) and AOT (Yang, Wei, and Yang 2021)), and salient object detection (iGAN (Mao et al. 2021) and LGVT (Zhang et al. 2021)). These results are collected from the AVS benchmark (Zhou et al. 2022), which are transferred from the original tasks to the AVS tasks.
@@
As shown in Table 1, our AVSegFormer exceeds these methods by large margins. For instance, on the S4 subset, AVSegFormer-R50 achieves an impressive mIoU of 76.45, which is 1.51 points higher than the best LGVT. Although LGVT has a better Swin-T (Liu et al. 2021) backbone, our AVSegFormer with ResNet-50 backbone still performs better regarding mIoU. In addition, AVSegFormer-PVTv2 produces an outstanding mIoU of 82.06 and an F-score of 89.9 on this subset, which is 7.12 mIoU and 2.6 F-score higher than LGVT, respectively.
On the MS3 subset, AVSegFormer-R50 outperforms the best iGAN with 6.64 mIoU and 8.4 F-score, while AVSegFormer-PVTv2 further raised the bar with an exceptional improvement of 15.47 mIoU and 14.9 F-score. On the AVSS subset, our AVSegFormer-R50 yields 24.93 mIoU and 29.3 F-score, and AVSegFormer-PVTv2 obtains an impressive performance of 36.66 mIoU and 42.0 F-score, surpassing AOT by 11.26 mIoU and 11.0 F-score, respectively.
@@
Comparison with AVSBench. Then, we compare our AVSegFormer with the AVSBench baseline, which is the current state-of-the-art method for audio-visual segmentation. As reported in Table 1, on the S4 subset, AVSegFormer-R50 achieves 3.66 mIoU and 1.1 F-score improvements over AVSBench-R50, while AVSegFormer-PVTv2 surpasses AVSBench-PVTv2 by 3.32 mIoU and 2.0 F-score. On the MS3 subset, AVSegFormer-PVTv2 surpasses AVSBench-PVTv2 with a margin of 1.65 mIoU and 5.0 F-score. On the AVSS subset, AVSegFormer-R50 and AVSegFormer-PVTv2 achieve significant results with an mIoU improvement of 4.75 and 6.89, and a substantial F-score improvement of 4.1 and 6.8. These results demonstrate that AVSegFormer outperforms the AVSBench baseline on all sub-tasks, becoming a new state-of-the-art method for audio-visual segmentation.
Ablation Study
@@
In this section, we conduct ablation experiments to verify the effectiveness of each key design in the proposed AVSegFormer. Specifically, we adopt PVTv2 (Wang et al. 2022b) as the backbone and conduct extensive experiments on the S4 and MS3 sub-tasks.
@@
Effect of learnable queries. We further investigated the impact of learnable queries in the decoder inputs. As shown in Table 2b, the improvement due to the learnable queries is relatively small in the single sound source task (S4), while it brings significant improvement in the multiple sound source task (MS3). This can be attributed to the complexity of sounding objects. We involve a more detailed discussion in the supplementary materials.
@@
Effect of audio-visual mixer. We then studied the impact of the audio-visual mixer on our model. Two versions are designed for this module, as illustrated in Figure 4. The cross-attention mixer (CRA) utilizes visual features as queries and audio features as keys/values for cross-attention, and the channel-attention mixer (CHA) introduced the mechanism of channel attention with audio features as queries and visual features as keys/values. As presented in Table 2c, the design of CHA brought greater performance improvement compared to CRA.
@@
In addition, we also visualize the mask feature before and after the audio-visual mixer, as shown in Figure 5. It is evident that for the sounding object (right girl and dog), the mixer effectively enhanced its features. Meanwhile, the non-sounding objects (left girl, table/bed, or background) experienced some degree of suppression. These findings align with our hypothesis and further substantiate the effectiveness of the audio-visual mixer.
Effect of auxiliary mixing loss. We finally conducted experiments to learn the impact of the auxiliary mixing loss. We train our model with and without the mixing loss, respectively, and report the testing results in Table 2d. It is demonstrated that the auxiliary mixing loss can help a lot in the final prediction.
@@
Qualitative analysis. We also present the visualization results of AVSegFormer compared with those of AVSBench on three audio-visual segmentation tasks in Figure 6. The visualization results clearly demonstrate that our method performs better. It has a strong ability in target localization and semantic understanding and can effectively identify the correct sound source and accurately segment the target object in multiple sound source scenes. These results highlight the effectiveness and robustness of our method.
Conclusion
In this paper, we propose AVSegFormer, a novel audio-visual segmentation framework that leverages the power of transformer architecture to achieve leading performance. Specifically, our method comprises four key components: (1) a transformer encoder building the mask feature, (2) a query generator providing audio-conditioned queries, (3) a dense audio-visual mixer dynamically adjusting interested visual features, and (4) a sparse audio-visual decoder separating audio sources and matching optimal visual features. These components provide a more robust audio-visual cross-modal representation, improving the AVS performance in different scenarios. Extensive experimental results demonstrate the superior performance of AVSegFormer compared to existing state-of-the-art methods.
Acknowledgement
This work is supported by the Natural Science Foundation of China under Grant 62372223.